### v1

# Hunting for extra dimensions in black hole shadows

Abstract
Observational data of the Sagittarius A* (Sgr A*) shadow released by the Event Horizon Telescope (EHT) are used to investigate eventual deviations in the black hole shadow radius, aiming to seek physics beyond the Standard Model (SM) coming from extra-dimensional theory. We consider the brane-world scenario described by the Randall-Sundrum model and determine the black hole shadow radius correction owing to the higher dimension. From data of the shadow radius in units of BH mass determined by KECK- and VLTI-based estimates, we imposed restrictions on the deviation obtained, and one sets an upper limit to the curvature radius of Anti-de Sitter ($\mathrm{AdS_{5}}$) spacetime $\ell\lesssim 4.3\times 10^{-2}\,\mathrm{AU}$ (at $95\%$ confidence level).
extra dimension, brane, black hole, quasinormal modes
pacs:
04.50.−h, 04.50.Gh, 11.10.Kk
I Introduction
Over the last century, attempts have been made to investigate the geometrical structure of the Universe from the proposal of the existence of extra dimensions (Nordstrom:1914ejq, ). Following the development of general relativity, a theoretical extra-dimensional model was built, which aimed to unify the electromagnetic and gravitational fields (Kaluza:1921tu, ; Klein:1926tv, ). Recently, the modern brane-world theories have renewed interest in higher dimensional theories (Antoniadis:1998ig, ; Arkani-Hamed:1998jmv, ; Arkani-Hamed:1998sfv, ; Randall:1999ee, ; Randall:1999vf, ). These proposed models arise in the unification schemes, where one expected that the unification of gravitational and gauge forces would occur near the $\mathrm{TeV}$ scale (Antoniadis:1998ig, ; Arkani-Hamed:1998jmv, ; Arkani-Hamed:1998sfv, ). In such case, the gauge interactions, and therefore the ordinary matter, get confined to a $4$-dimensional hypersurface ($3$-brane), whereas gravity lives in a higher-dimensional spacetime known as bulk (Randall:1999vf, ; Randall:1999ee, ). Thus, gravity would be diluted in the higher-dimensional space, spreading throughout the bulk, explaining, therefore, the observed weakness of gravity compared to other fundamental interactions of nature (Antoniadis:1998ig, ; Arkani-Hamed:1998jmv, ; Arkani-Hamed:1998sfv, ; Randall:1999ee, ; Randall:1999vf, ). Consequently, this is a promising scenario to address the hierarchy problem that deals with the discrepancy between the gravitational and electroweak fundamental energy scales without relying on low-energy supersymmetry or technicolor (Antoniadis:1998ig, ; Arkani-Hamed:1998jmv, ; Arkani-Hamed:1998sfv, ).
In this framework, although the Standard Model (SM) fields are trapped on our 4-dimensional manifold, events of high-energy collisions can, for instance, eventually cause the SM particles to be kicked off into the new dimensions (ParticleDataGroup:2022pth, ). Thus, collider experiments have been employed to seek for evidence of large extra dimensions in Arkani-Hamed-Dimopoulos-Dvali (ADD) model, and it was found that the most stringent limits on the Planck mass of higher-dimensional spacetime come from ATLAS (CMS) missing energy experiment, giving $M_{D}>5.9\,\text{–}\,11.2\,\mathrm{TeV}$ (ATLAS:2021kxv, ). In its turn, analysis from astrophysical and cosmological origin, such as light KK gravitons production in stars, has also provided the strongest bounds so that $M_{D}>1700\,(76)\,\mathrm{TeV}$ for $\delta=2\left(3\right)$ (Hannestad:2003yd, ). Further discussions have attempted to seek traces of extra dimensions in several physical systems, such as in rotating torsion balance (Adelberger:2009zz, ; Murata:2014nra, ) – wherein one tests Newton’s law at sub-mm distances –, spectroscopy and atomic physics (Dahia:2015xxa, ; Dahia:2015bza, ; Dahia:2017jya, ; Lemos:2019rhj, ; Lemos:2021xpi, ; Luo:2006ad, ; Zhi-gang:2007swh, ; Salumbides:2015qwa, ), and black hole production (CMS:2018ozv, ; ATLAS:2015yln, ).
On the other hand, the model proposed by Randall and Sundrum (RS) (Randall:1999ee, ; Randall:1999vf, ) describes our universe by postulating it as a five-dimensional Anti-de Sitter ($\mathrm{AdS_{5}}$) spacetime with a warped geometry. In this scenario, bounds on the $\mathrm{AdS_{5}}$ curvature radius $\ell$ have also been imposed from astrophysical observations. For instance, the investigation of the persistence of X-ray binaries containing a black hole in an $\mathrm{AdS_{5}}$ brane-world shows that $\ell\lesssim 10^{-2}\,\mathrm{mm}$ (Emparan:2002jp, ). Furthermore, recent work has also aimed to search for signals of the extra dimension on events of black hole gravitational waves (Bohra:2023vls, ; Chrysostomou:2022evl, ). In such a way, the study of Hawking radiation emitted by stellar-mass black holes found that $\ell\lesssim 5\,\mathrm{\mu m}$ from gravitational wave measurements (McWilliams:2009ym, ), while Ref. (Vagnozzi:2019apd, ) – analyzing the shadow of M87* – imposes constraints on the $\mathrm{AdS_{5}}$ radius such that $\ell\lesssim 170\,\mathrm{AU}$. The multi-messenger cosmology, which considers joint observations of gravitational waves and electromagnetic counterparts, has been used to set limits on the radius of curvature of the higher-dimensional spacetime, $\ell\lesssim 0.535\,\mathrm{Mpc}$ (Visinelli:2017bny, ). On its turn, Ref. (Mishra:2021waw, ) constrains the tidal charge of a rotating brane-world black hole.
It is known that the black holes only recently have been experimentally probed. Such observations may provide us with significant insights into the structure of spacetime. In this context, the LIGO-VIRGO collaboration has been responsible for presenting empirical data on gravitational wave events from the merger of binary black hole system and neutron stars (LIGOScientific:2016aoc, ; LIGOScientific:2017vwq, ). In the black hole merger process a ringdown phase arises from the gravitational wave signal, and in this scenario the study of quasinormal modes (QNMs) has been attracting interest in recent years (cardoso2003quasinormal, ; cardoso2019parametrized, ; cruz2020polar, ; anacleto2021quasinormal, ; campos2022quasinormal, ). Quasinormal modes are solutions of perturbation equations that have complex characteristic oscillation frequencies that satisfy specific boundary conditions (konoplya2011quasinormal, ). The QNMs can be obtained using approximative methods, such as Wentzel-Kramers-Brillouin (WKB) approximation or other numerical methods, in different orders of precision (schutz1985black, ; iyer1987black, ; konoplya2003quasinormal, ; konoplya2019higher, ).
In higher-dimensional theories, QNMs have gained increasing attention in recent years (Chakraborty:2017qve, ; Dey:2020pth, ; Hashemi:2019jlt, ; konoplya2003quasinormal, ; Cardoso:2003vt, ; Anacleto:2023ntm, ), once one expects such investigations may reveal signals of extra dimensions. In turn, recent measurements of the Event Horizon Telescope (EHT) provided images of supermassive black hole shadow at the M87* galaxy center (EventHorizonTelescope:2019dse, ; EventHorizonTelescope:2019ggy, ), in addition to the shadow of compact source Sagittarius A* (Sgr A*), a supermassive black hole in the center of the Milky Way (EventHorizonTelescope:2022xqj, ; EventHorizonTelescope:2022wkp, ). It is noteworthy that, in this scenario, the search for physics beyond the SM has been based on analysis of the shadow radius and deviation from the circularity $\Delta C$ of the shadow. However, whereas the circularity has been quoted for the M87* measurements, such information for Sgr A* observational data is not available. Therefore, the approach employed in Ref. (Vagnozzi:2019apd, ) cannot be applied to test fundamental physics from the circularity deviations of Sgr A*, and one must recover to investigations based on its shadow size. In the present work, we aim to determine conservative constraints on the $\mathrm{AdS_{5}}$ radius of curvature from observational data of the shadow of Sgr A*. For this purpose, we will consider a black hole in the RS background in the weak field regime, described by the Garriga-Tanaka metric (Garriga:1999yh, ). This solution has been obtained by assuming a classic approach that provides the linearized brane-world gravity. As we will see, our analysis allowed us to place the bounds on the extra-dimensional curvature radius $\ell$ in an alternative way, avoiding the issue of lack of circularity measures for Sgr A* and exploring its inaccurate spin measurements (Banerjee:2022jog, ; Fragione:2020khu, ). These findings are supported by the WKB shadow estimate obtained from the study of quasinormal modes in the eikonal limit.
The organization of this paper is as follows: In section II, we discuss the propagation of a massless scalar field in the background of a black hole on a thin brane-world known as the Randall-Sundrum model (RSII), described by a warped geometry theory. In Sec. III, we have determined the quasinormal modes for a scalar perturbation on this background wherein eventual effects of the fifth dimension must become relevant. Therefore, we use the WKB approximation, aiming to determine the quasinormal frequencies in the eikonal limit. In Sec. IV, we investigate the null geodesic curves to find the shadow radius of the RSII black hole, and we compare it with the prediction obtained from the eikonal quasinormal frequencies. In this case, we find deviations in the shadow radius of the black hole owing to the extra-dimensional curvature radius. From the EHT observational data, we present the imposed constraints on the $\mathrm{AdS_{5}}$ curvature radius $\ell$. The concluding remarks and the main results are given in Sec. V.
II Black holes on a thin brane: the radial equation
In this section, we aim to describe the dynamics of a massless scalar field propagating on a thin brane-world background. Thus, we start by considering the Garriga-Tanaka solution that, at the first-order approximation of $GM$, describes the metric of a matter distribution with mass $M$ localized in the brane (Garriga:1999yh, ),
$ds^{2}=-f(r)dt^{2}+\dfrac{dr^{2}}{g(r)}+r^{2}d\Omega^{2},$ (1)
where, in the weak field limit $\left(r\gg\ell\right)$ from the horizon of matter after the collapse, we get
$f(r)=1-\dfrac{2M}{r}-\dfrac{4M\ell^{2}}{3r^{3}},\qquad\text{and}\qquad g(r)=1-\dfrac{2M}{r}-\dfrac{2M\ell^{2}}{3r^{3}},$ (2)
in which $\ell$ is the curvature radius of $\mathrm{AdS_{5}}$, and we consider unities $c=G=1$. In this scenario, we may discuss the propagation of a massless scalar field and obtain the radial equation by using the Klein–Gordon equation
$\frac{1}{\sqrt{-g}}\partial_{\mu}\left(\sqrt{-g}g^{\mu\nu}\partial_{\nu}\Psi\right)=0.$ (3)
For the sake of simplicity, one may use the following separation of variables
$\Psi_{\omega lm}\left(\vec{r},t\right)=\frac{\mathcal{R}_{\omega l}(r)}{r}Y_{l\,m}(\theta,\phi)e^{-i\omega t},$ (4)
where $Y_{l\,m}(\theta,\phi)$ are the spherical harmonics, $\omega$ is the frequency and $l$ is the multipole number. Thus, substituting the expression (4) into equation (3), one finds the following radial equation for $\mathcal{R}_{\omega l}$
$\Lambda(r)\frac{d}{dr}\left[\Lambda(r)\frac{d\mathcal{R}_{\omega l}}{dr}\right]+\left(\omega^{2}-V_{eff}\right)\mathcal{R}_{wl}=0.$ (5)
Here, the effective potential is given by
$V_{eff}=\dfrac{\Lambda(r)}{r}\dfrac{d\Lambda(r)}{dr}+\dfrac{f(r)}{r^{2}}l(l+1),$ (6)
where $\Lambda(r)=\sqrt{g(r)f(r)}$. In order to analyze the radial equation (5), by considering the asymptotic limits, we can introduce a new coordinate $dx=\Lambda(r)^{-1}dr$ (commonly called tortoise coordinate) so that equation (5) reduces to a Schrödinger-type equation, as follows
$\dfrac{d^{2}\mathcal{R}_{\omega l}}{dx^{2}}+\left(\omega^{2}-V_{eff}\right)\mathcal{R}_{\omega l}=0.$ (7)
As we will see below, this analysis will allow us to estimate the shadow radius by calculating the quasinormal modes in the eikonal limit, which corresponds to the limit to a large multipole number $l$.
III Eikonal quasinormal modes in the warped brane-world
Quasinormal modes are solutions of the perturbation equation (7) that satisfy specific boundary conditions. In this case, it is possible to consider two different solutions obtained to the radial equation written in terms of the new coordinate by assuming some set of boundary conditions (berti2009quasinormal, ). Furthermore, these solutions present a discrete quasinormal frequency spectrum $\omega=\omega_{R}+i\omega_{I}$, where the real part ($\omega_{R}$) determines the oscillation frequency, while the imaginary part ($\omega_{I}$) determines the decay rate. In a regime for which the real part of the QNM frequency is large, one has the called eikonal limit. On its turn, WKB approximation has been applied to obtain the quasinormal modes in black hole for decades (schutz1985black, ; iyer1987black, ; konoplya2003quasinormal, ). This method is widely employed in the literature due to its practical means of finding approximating solutions compared to numerical methods, in addition to providing accurate predictive results (konoplya2019higher, ; anacleto2021quasinormal, ; campos2022quasinormal, ). The WKB method in the eikonal limit yields an accurate estimate for the quasinormal frequencies from the wave equation (7). Specifically, the QNM frequencies in this particular limit can be obtained from the following expression
$\dfrac{i\left(\omega_{n}^{2}-V_{0}\right)}{\sqrt{-2V^{\prime\prime}_{0}}}=n+\dfrac{1}{2},$ (8)
where $n$ is the harmonic number, $V_{0}$ is the maximum effective potential evaluated at the point $r_{0}$, the prime () denotes derivation with respect to the tortoise coordinate $x$. Furthermore, the effective potential (6) at the eikonal limit $(l\rightarrow\infty)$ is $V_{eff}\approx f(r)l^{2}/r^{2}$, so that the equation (8) can be written in the form
$\omega_{n}\approx\dfrac{l\sqrt{f(r_{0})}}{r_{0}}-\dfrac{i(n+1/2)}{\sqrt{2}}\sqrt{-\dfrac{r_{0}^{2}}{f(r_{0})}\dfrac{d^{2}}{dx^{2}}\left[\dfrac{f(r)}{r^{2}}\right]_{r=r_{0}}},$ (9)
which corresponds to the quasinormal frequency to the overtone mode $n$. In Fig. 1, one can see that both parts, real and imaginary, of eikonal quasinormal frequencies remain finite. In principle, the value of the $\mathrm{AdS_{5}}$ curvature radius $\ell$ must be constrained just from the empirical data such that the quasinormal frequencies will tend to zero in the eikonal limit for large values of the $\ell$-parameter. Furthermore, the result (9) obtained by assuming the eikonal limit will be compared with the null geodesics parameters, as we will see in the next section. From computing the deviations on the black hole shadow radius in the brane-world model, we will impose bounds on the curvature radius $\mathrm{AdS_{5}}$.
Figure 1: Real and negative imaginary parts of the eikonal quasinormal frequencies as function of the $\mathrm{AdS_{5}}$ curvature radius.
IV Corrected Shadow Radius of the Brane-world Black Hole
In this section, we will study the null geodesic, aiming to investigate the influence of the extra dimension on the critical impact and critical radius parameters. This approach will also allow us to determine the radius of the shadow of the Garriga-Tanaka (RSII) black hole. From this investigation, we will find bounds on the curvature radius of the $\mathrm{AdS_{5}}$ spacetime. We begin by considering the following Lagrangian density
$\mathcal{L}\equiv\frac{1}{2}g_{\mu\nu}\dot{x}^{\mu}\dot{x}^{\nu},$ (10)
where the dot denotes the derivative with respect to the affine parameter. Now, applying the line element (1), we have
$2\mathcal{L}=f(r)\dot{t}^{2}-\dfrac{\dot{r}^{2}}{g(r)}-r^{2}\left(\dot{\theta}^{2}+\sin^{2}\theta\dot{\phi}^{2}\right).$ (11)
For the sake of simplicity, at this point, by setting the angle $\theta=\pi/2$, one can examine the motion in an equatorial plane. Thus, we can find the equation of motion for the Hamilton-Jacobi equation
$E=f(r)\dot{t},\qquad\quad L=r^{2}\dot{\phi},$ (12)
where $E$ and $L$ are the constants corresponding to energy and angular momentum, respectively. For a null geodesic, we have $g_{\mu\nu}\dot{x}^{\mu}\dot{x}^{\nu}=0$, and we can find the following expression
$\dot{r}^{2}+\mathcal{V}(r)=0,\qquad\mathcal{V}(r)=\dfrac{g(r)}{f(r)}E^{2}-g(r)\dfrac{L^{2}}{r^{2}}.$ (13)
Now, just for convenience, we can introduce a new variable $u=1/r$ to write the orbital equation
$\left(\dfrac{du}{d\phi}\right)^{2}=\dfrac{\left(3-6Mu-2M\ell^{2}u^{3}\right)}{\left(3-6Mu-4M\ell^{2}u^{3}\right)b^{2}}-\left(1-2Mu-\dfrac{2M\ell^{2}u^{3}}{3}\right)u^{2},$ (14)
$\dfrac{d^{2}u}{d\phi^{2}}=\dfrac{2M\ell^{2}\left(3-4Mu\right)u^{2}}{b^{2}\left(6Mu+4M\ell^{2}u^{3}-3\right)^{2}}-\left(1-3Mu-\dfrac{5M\ell^{2}u^{3}}{3}\right)u.$ (15)
Here, we have that $b=L/E$ is the impact parameter defined as the perpendicular distance (measured from infinity) between the geodesic and the parallel line that passes through the origin. Using the following critical conditions, (i) $\mathcal{V}(r_{c})=0$ and (ii) $d\mathcal{V}(r)/dr\big{|}_{r=r_{c}}=0$, we can obtain the critical impact parameter from a system of equations with which one relates the impact parameter to the critical radius, as follows
$b_{c}^{2}=\dfrac{r_{c}^{2}}{f(r_{c})}=\dfrac{3r_{c}^{5}}{3r_{c}^{3}-6Mr_{c}^{2}-4M\ell^{2}}.$ (16)
Notice that $r_{c}$ corresponds to the radius of photon circular orbit. On its turn, the critical conditions expressed by (ii) yield a polynomial equation that is used to determine the critical radius:
$0\displaystyle 20\ell^{4}M^{2}+6\ell^{2}M\left(13M-6r_{c}\right)r_{c}^{2}+9r_{c}^{4}\left(6M^{2}-5Mr_{c}+r_{c}^{2}\right)=$ 0 . (17)
We must highlight that, here, we want to study the trajectory of a ray of light in the spherically symmetric Garriga-Tanaka background. Specifically, when we analyze the ray in a plane, we observe that any ray of light that starts at a particular angle $\theta$ must keep the same angle throughout its path. In Figure 2, we have found some results for the critical impact parameter from the numerical solution of the orbit equations (14) and (15). In this case, we have obtained the geodesic trajectory (blue thick curve) computed for the impact parameter $b=5.5M$. Therefore, as a direct influence of the existence of the extra-dimensional parameter $\ell$, we observed an increase in the absorption section. Furthermore, in Figure 2, whereas the black disk depicts the limit of the event horizon, the critical radius for the photon sphere is shown as an internal dotted circle. In turn, the critical impact parameter (black hole shadow) is shown by an external dashed circle.
Figure 2: Null geodesic in polar coordinates for different values of $\ell$, in which we consider a test beam of light by keeping the impact parameter, $b=5.5M$, fixed.
Finally, one can estimate the size of the radius of the black hole shadow by using the study of geodesic curves that we might express via celestial coordinates (vazquez2003strong, )
$\alpha$ $=$ $\lim\limits_{r_{o}\to\infty}\left[-r_{o}^{2}\sin\theta_{o}\dfrac{d\phi}{dr}\Bigr{\rvert}_{\theta=\theta_{o}}\right],$ (18)
$\beta$ $=$ $\lim\limits_{r_{o}\to\infty}\left[r_{o}^{2}\dfrac{d\theta}{dr}\Bigr{\rvert}_{\theta=\theta_{o}}\right],$ (19)
where $\left(r_{o},\theta_{o}\right)$ is the position of the observer at infinity. Thus, for an observer on an equatorial plane $\theta_{o}=\pi/2$, we have the following relationship for the shadow radius
$R_{s}\equiv\sqrt{\alpha^{2}+\beta^{2}}=\sqrt{\dfrac{3r_{c}^{5}}{3r_{c}^{3}-6Mr_{c}^{2}-4M\ell^{2}}}.$ (20)
Let us now expand the radius of the shadow up to order $\left(\ell^{2}/M^{2}\right)$ by assuming that the contribution arising from extra-dimension is a small correction to $R_{s}$, $\ell/M\ll 1$, so that
$R_{s}\approx 3\sqrt{3}M\left[1+\frac{2}{27}\frac{\ell^{2}}{M^{2}}+\mathcal{O}\left(\frac{\ell^{4}}{M^{4}}\right)\right].$ (21)
On the other hand, for a spherically symmetric and asymptotically flat static geometry, Cardoso et al. (cardoso2009geodesic, ) showed that, in the eikonal limit, the quasinormal modes are determined by the parameters of the null circular geodesic so that $\Omega_{c}=\dot{\phi}/\dot{t}$ is the angular velocity for the last null circular orbit, and $\lambda=\sqrt{\mathcal{V}^{\prime\prime}(r)/2\dot{t}^{2}}$ is the Lyapunov exponent that determines the unstable time scale of the orbit. By directly comparing with equation (9), we can write the equation for quasinormal frequencies in the form
$\omega_{n}=\Omega_{c}l-i(n+1/2)|\lambda|.$ (22)
It is worth to mention that there are violations in this relationship, as shown by Konoplya and Stuchlík (konoplya2017eikonal, ) when considering a black hole in the Einstein-Loverlock theory. Since we know the form of the angular velocity $\Omega_{c}$, one can use equations (12) and (16) to obtain its relationship with the critical impact parameter $\Omega_{c}=\left(f(r_{c})b_{c}\right)/r_{c}^{2}=1/b_{c}$. This equation highlights the correspondence relationship between the maximum value point of the effective potential ($r_{0}$) and the radius of the photon circular orbit ($r_{c}$) [see equations (9) and (22)]. Finally, in the eikonal regime $l\gg 1$, we have the relationship between the real part of the quasinormal modes and the shadow radius,
$Re(\omega)=\lim_{l\gg 1}\left(\dfrac{l}{R_{s}}\right).$ (23)
Notice that the estimate (23) may be employed to confirm the prediction (21). Hence, we can use it to compare with experimental data in order to constrain the found deviations. Besides, equation (22) shows that the real and imaginary parts of the quasinormal modes in the eikonal limit are related to the orbital angular velocity and the instability time scale of circular null geodesics, respectively. The instability of null circular orbits can be determined by the instability exponent, which is defined by the ratio of the Lyapunov exponent to the angular frequency. In Fig.3, we see the variation of the instability exponent in relation to the radius $r_{c}$. The maximum of the instability increases with the increase in the contribution coming from the extra dimension. On the other hand, the instability decreases as $r_{c}$ increases.
Figure 3: Lyapunov exponent normalized by the angular velocity $\left(\lambda/\Omega_{c}\right)$, as function of $r_{c}/M$ for some values of $\ell$.
IV.1 Constraints on the $\mathrm{AdS_{5}}$ curvature radius
As we saw above, we obtained deviations in the black hole shadow radius owing to the existence of an extra dimension. On its turn, if we turn off the extra-dimensional effects ($\ell=0$), we recover the well-known results for the Schwarzschild background. Although this independent approach allows us to search for traces of extra dimensions, imposing constraints on deviations found, once the curvature radius $\ell$ is positively correlated to the shadow radius, and given the low-precision measurements, thus we must have a weakening in the predicted bounds compared to sub-mm gravity experiments, for instance.
In our analysis, we will follow the methodology developed in (vagnozzi2022horizon, ), aiming to obtain bounds on the $\mathrm{AdS_{5}}$ curvature radius. Therefore, the observed angular radius – which has ring-like features in the horizon-scale EHT image of Sgr A* –, must be compared to the theoretically determined angular radius of the black hole shadow. From the precise determination of both mass and distance of Sgr A*, the central supermassive black hole of our galaxy, we can verify the accordance, within the uncertainty allowed, between the predicted theoretical results by this study and the empirical data found by the EHT observations (akiyama2022millimeter, ). On its turn, the mass and distance of Sgr A* have been studied over the last few decades by exploring the dynamics of star clusters and, in particular, the movement of stars close to the center of the galaxy. According to Ref. (vagnozzi2022horizon, ), the shadow radius measurements are restricted to the following regions determined by the average between the estimates based on the VLTI and Keck instruments, within $1\sigma$ and $2\sigma$ intervals, respectively,
$4.55\lesssim R_{s}/M\lesssim 5.22,\qquad\text{and}\qquad 4.21\lesssim R_{s}/M\lesssim 5.56$ (24)
In Figure 4a, we show the expected deviations in the behavior of the shadow radius owing to the extra dimension. As we see, the increase in the $\mathrm{AdS_{5}}$ curvature radius $\ell$ results in a consequent increase in the radius of the black hole shadow. In our analysis, we have considered the average in the mass from KECK- and VLTI-based estimates. Therefore, from the empirical data obtained by observations of EHT, we can impose the limits on curvature radius at the $95\%$ confidence level (CL), such that $\ell\lesssim 6.5\times 10^{9}\,\mathrm{m}\simeq 4.3\times 10^{-2}\,\mathrm{AU}$, as shown in Figure 4b. At this point, we can compare our results with another which investigates similar physical systems. In this case, the absence of deviations on observation of the highly circular shadow of black hole M87* provides constraints on the $\mathrm{AdS_{5}}$ radius so that $\ell\lesssim 170\,\mathrm{AU}$ (Vagnozzi:2019apd, ). In turn, the first bounds obtained on the extra dimensions by the study of the time-lag between the detection of gravitational waves (GW) and electromagnetic (EM) signals from the merging of a binary system detected by the event GW170817 have established at $68\%$ CL an upper limit of $\ell\lesssim 0.535\,\mathrm{Mpc}\simeq 1.10\times 10^{11}\,\mathrm{AU}$ (Visinelli:2017bny, ), whereas our analysis found $\ell\lesssim 1.0\times 10^{-2}\,\mathrm{AU}$ for the same confidence interval. It is worth mentioning that, as we have seen, our study yields stronger constraints than those found by similar physical systems, according to the best of our knowledge. Ultimately, although we obtained weaker constraints compared to results obtained from sub-mm scale gravity experiments, this approach has set independent limits and can aid the investigation of the geometric structure of the universe.
Figure 4: The left panel shows the changes in the behavior of the shadow radius with the variation of the $\mathrm{AdS_{5}}$ radius $\ell$. In turn, in the right panel, we found bounds from the direct comparison between the results obtained by our theoretical analysis with the EHT experimental data.
V Final Remarks
Recently, higher-dimensional theories have drawn increasing attention. As has been pointed out, the seek for evidence of large extra dimensions has been performed in several experimental researches, making its detection a challenging problem in itself. In this work, we have studied the quasinormal modes of a black hole in a brane-world model, aiming to search for traces of extra dimensions. By employing the WKB method, we compute the eikonal quasinormal frequencies. This complementary analysis also allowed us to confirm the theoretical prediction for the black hole shadow radius from the convergence between the WKB- and null geodesics-method estimates.
From recent measurements of the shadow image of Sgr A* – a supermassive black hole in the center of our galaxy – we estimate deviations owing to an extra dimension in the RSII scenario on the theoretical shadow radius predicted. Given the inaccuracy involving the measured spin of Sgr A*, which may even be zero but still can provide compatible results with recent measures (Fragione:2020khu, ), we explore the possibility of describing them as a non-rotating RSII black hole. In this case, we have considered the Garriga-Tanaka solution that takes a classical approach to linearized brane-world gravity and describes a black hole solution in the weak field limit.
As we saw, the curvature radius of the $\mathrm{AdS_{5}}$ spacetime is positively correlated to the shadow radius so that the increasing/decreasing in one leads to correspondingly increasing/decreasing in the other. Furthermore, we found that $\ell=0$, i.e., in the absence of new physics coming from extra dimension, one recovers the standard estimates expected for the Schwarzschild background. By investigating the null geodesic curves, we determine the shadow of the RSII black hole. From recent observational data of the shadow of Sgr A* released by the Event Horizon Telescope collaboration, one may impose constraints on the curvature radius of $\mathrm{AdS_{5}}$ spacetime. Therefore, we find more stringent limits on $\ell\lesssim 4.3\times 10^{-2}\,\mathrm{AU}$ (at $95\%$ CL) than previously obtained for similar physical systems, to the best of our knowledge. Moreover, non-zero predicted curvature radius also leads to agreement with the observed shadow radius within $68\%$ confidence interval, providing $\ell\lesssim 1.0\times 10^{-2}\,\mathrm{AU}$.
We would like to highlight that although our analysis provides weaker bounds than those obtained from collider experiments or sub-mm precision gravity tests, which have set limits on the non-Newtonian force, in this study, we have provided conservative bounds since limits ourselves to investigate the shadow radius in a non-rotating RSII-type black hole background. Further, the bounds reported by us have been obtained in a manner complementary to the studies previously mentioned. Finally, the newly begun era of observational astronomy opens several avenues for future research that might be explored to obtain new constraints on the $\mathrm{AdS_{5}}$ curvature radius, for instance, or even to test extensions of general relativity, therefore shedding light into seeking for physics beyond the SM.
Acknowledgements.
We would like to thank CNPq, CAPES and CNPq/PRONEX/FAPESQ-PB (Grant No. $165/2018$), for partial financial support. FAB acknowledges support from CNPq (Grant No. $309092/2022-1$). ASL acknowledges support from CAPES (Grant No. $88887.800922/2023-00$). JAVC would like to thank FAPESQ-PB/CNPq (Grant No. $077/2022$) for financial support.

### v2

@@
 Abstract
-Observational data of the Sagittarius A* (Sgr A*) shadow released by the Event Horizon Telescope (EHT) are used to investigate eventual deviations in the black hole shadow radius, aiming to seek physics beyond the Standard Model (SM) coming from extra-dimensional theory.
-We consider the brane-world scenario described by the Randall-Sundrum model and determine the black hole shadow radius correction owing to the higher dimension.
-From data of the shadow radius in units of BH mass determined by KECK- and VLTI-based estimates, we imposed restrictions on the deviation obtained, and one sets an upper limit to the curvature radius of Anti-de Sitter ($\mathrm{AdS_{5}}$) spacetime $\ell\lesssim 4.3\times 10^{-2}\,\mathrm{AU}$ (at $95\%$ confidence level).
+Observational data of the Sagittarius A* (Sgr A*) shadow released by the Event Horizon Telescope are used to investigate eventual deviations in the black hole shadow radius, aiming to seek physics beyond the Standard Model coming from extra-dimensional theory.
+We consider the braneworld scenario described by the Randall-Sundrum model and determine the black hole shadow radius correction owing to the higher dimension.
+From data of the shadow radius in units of black hole mass determined by Keck- and Very Large Telescope Interferometer (VLTI)-based estimates, we imposed restrictions on the deviation obtained, and one sets an upper limit to the curvature radius of anti-de Sitter ($\mathrm{AdS_{5}}$) spacetime $\ell_{\left(\text{Keck}\right)}\lesssim 3.8\times 10^{-2}\,\mathrm{AU}$ and $\ell_{\left(\text{VLTI}\right)}\lesssim 1.6\times 10^{-2}\,\mathrm{AU}$ (at $68\%$ confidence level), respectively.
 extra dimension, brane, black hole, quasinormal modes
@@
 I Introduction
-Over the last century, attempts have been made to investigate the geometrical structure of the Universe from the proposal of the existence of extra dimensions (Nordstrom:1914ejq, ).
-Following the development of general relativity, a theoretical extra-dimensional model was built, which aimed to unify the electromagnetic and gravitational fields (Kaluza:1921tu, ; Klein:1926tv, ).
-Recently, the modern brane-world theories have renewed interest in higher dimensional theories (Antoniadis:1998ig, ; Arkani-Hamed:1998jmv, ; Arkani-Hamed:1998sfv, ; Randall:1999ee, ; Randall:1999vf, ).
+Over the past century, attempts have been made to investigate the geometrical structure of the Universe from the proposal of the existence of extra dimensions (Nordstrom:1914ejq, ).
+Following the development of general relativity (GR), a theoretical extra-dimensional model was built, which aimed to unify the electromagnetic and gravitational fields (Kaluza:1921tu, ; Klein:1926tv, ).
+Recently, the modern braneworld theories have renewed interest in higher dimensional theories (Antoniadis:1998ig, ; Arkani-Hamed:1998jmv, ; Arkani-Hamed:1998sfv, ; Randall:1999ee, ; Randall:1999vf, ).
 These proposed models arise in the unification schemes, where one expected that the unification of gravitational and gauge forces would occur near the $\mathrm{TeV}$ scale (Antoniadis:1998ig, ; Arkani-Hamed:1998jmv, ; Arkani-Hamed:1998sfv, ).
-In such case, the gauge interactions, and therefore the ordinary matter, get confined to a $4$-dimensional hypersurface ($3$-brane), whereas gravity lives in a higher-dimensional spacetime known as bulk (Randall:1999vf, ; Randall:1999ee, ).
+In such a case, the gauge interactions, and therefore the ordinary matter, get confined to a four-dimensional hypersurface (three-brane), whereas gravity lives in a higher-dimensional spacetime known as bulk (Randall:1999vf, ; Randall:1999ee, ).
 Thus, gravity would be diluted in the higher-dimensional space, spreading throughout the bulk, explaining, therefore, the observed weakness of gravity compared to other fundamental interactions of nature (Antoniadis:1998ig, ; Arkani-Hamed:1998jmv, ; Arkani-Hamed:1998sfv, ; Randall:1999ee, ; Randall:1999vf, ).
 Consequently, this is a promising scenario to address the hierarchy problem that deals with the discrepancy between the gravitational and electroweak fundamental energy scales without relying on low-energy supersymmetry or technicolor (Antoniadis:1998ig, ; Arkani-Hamed:1998jmv, ; Arkani-Hamed:1998sfv, ).
-In this framework, although the Standard Model (SM) fields are trapped on our 4-dimensional manifold, events of high-energy collisions can, for instance, eventually cause the SM particles to be kicked off into the new dimensions (ParticleDataGroup:2022pth, ).
-Thus, collider experiments have been employed to seek for evidence of large extra dimensions in Arkani-Hamed-Dimopoulos-Dvali (ADD) model, and it was found that the most stringent limits on the Planck mass of higher-dimensional spacetime come from ATLAS (CMS) missing energy experiment, giving $M_{D}>5.9\,\text{–}\,11.2\,\mathrm{TeV}$ (ATLAS:2021kxv, ).
-In its turn, analysis from astrophysical and cosmological origin, such as light KK gravitons production in stars, has also provided the strongest bounds so that $M_{D}>1700\,(76)\,\mathrm{TeV}$ for $\delta=2\left(3\right)$ (Hannestad:2003yd, ).
-Further discussions have attempted to seek traces of extra dimensions in several physical systems, such as in rotating torsion balance (Adelberger:2009zz, ; Murata:2014nra, ) – wherein one tests Newton’s law at sub-mm distances –, spectroscopy and atomic physics (Dahia:2015xxa, ; Dahia:2015bza, ; Dahia:2017jya, ; Lemos:2019rhj, ; Lemos:2021xpi, ; Luo:2006ad, ; Zhi-gang:2007swh, ; Salumbides:2015qwa, ), and black hole production (CMS:2018ozv, ; ATLAS:2015yln, ).
+In this framework, although the Standard Model (SM) fields are trapped on our four-dimensional manifold, events of high-energy collisions can, for instance, eventually cause the SM particles to be kicked off into the new dimensions (ParticleDataGroup:2022pth, ).
+Thus, collider experiments have been employed to seek evidence of large extra dimensions in the Arkani-Hamed-Dimopoulos-Dvali (ADD) model, and it was found that the most stringent limits on the Planck mass of higher-dimensional spacetime come from the ATLAS (CMS) missing energy experiment, giving $M_{D}>5.9–11.2\,\mathrm{TeV}$ (ATLAS:2021kxv, ).
+In its turn, analysis from astrophysical and cosmological origin, such as light Kaluza-Klein (KK) graviton production in stars, has also provided the strongest bounds so that $M_{D}>1700\,(76)\,\mathrm{TeV}$ for $\delta=2\left(3\right)$ (Hannestad:2003yd, ).
+Further discussions have attempted to seek traces of extra dimensions in several physical systems, such as in rotating torsion balance (Adelberger:2009zz, ; Murata:2014nra, )—wherein one tests Newton’s law at sub-mm distances—spectroscopy and atomic physics (Dahia:2015xxa, ; Dahia:2015bza, ; Dahia:2017jya, ; Lemos:2019rhj, ; Lemos:2021xpi, ; Luo:2006ad, ; Zhi-gang:2007swh, ; Salumbides:2015qwa, ; Lemos:2022vhw, ), and black hole production (CMS:2018ozv, ; ATLAS:2015yln, ).
 On the other hand, the model proposed by Randall and Sundrum (RS) (Randall:1999ee, ; Randall:1999vf, ) describes our Universe by postulating it as a five-dimensional anti-de Sitter ($\mathrm{AdS_{5}}$) spacetime with a warped geometry.
 In this scenario, bounds on the $\mathrm{AdS_{5}}$ curvature radius $\ell$ have also been imposed from astrophysical observations.
-For instance, the investigation of the persistence of X-ray binaries containing a black hole in an $\mathrm{AdS_{5}}$ brane-world shows that $\ell\lesssim 10^{-2}\,\mathrm{mm}$ (Emparan:2002jp, ).
+For instance, the investigation of the persistence of x-ray binaries containing a black hole in an $\mathrm{AdS_{5}}$ braneworld shows that $\ell\lesssim 10^{-2}\,\mathrm{mm}$ (Emparan:2002jp, ).
 Furthermore, recent work has also aimed to search for signals of the extra dimension on events of black hole gravitational waves (Bohra:2023vls, ; Chrysostomou:2022evl, ).
 In such a way, the study of Hawking radiation emitted by stellar-mass black holes found that $\ell\lesssim 5\,\mathrm{\mu m}$ from gravitational wave measurements (McWilliams:2009ym, ), while Ref.
-(Vagnozzi:2019apd, ) – analyzing the shadow of M87* – imposes constraints on the $\mathrm{AdS_{5}}$ radius such that $\ell\lesssim 170\,\mathrm{AU}$.
-The multi-messenger cosmology, which considers joint observations of gravitational waves and electromagnetic counterparts, has been used to set limits on the radius of curvature of the higher-dimensional spacetime, $\ell\lesssim 0.535\,\mathrm{Mpc}$ (Visinelli:2017bny, ).
+(Vagnozzi:2019apd, ) —analyzing the shadow of M87*—imposes constraints on the $\mathrm{AdS_{5}}$ radius such that $\ell\lesssim 170\,\mathrm{AU}$.
+The multimessenger cosmology, which considers joint observations of gravitational waves and electromagnetic counterparts, has been used to set limits on the radius of curvature of the higher-dimensional spacetime, $\ell\lesssim 0.535\,\mathrm{Mpc}$ (Visinelli:2017bny, ).
 On its turn, Ref.
-(Mishra:2021waw, ) constrains the tidal charge of a rotating brane-world black hole.
+(Mishra:2021waw, ) constrains the tidal charge of a rotating braneworld black hole, whereas in Ref.
+(Hou2021, ) one investigates how the extra dimension affects the shape of the black hole shadows.
 It is known that the black holes only recently have been experimentally probed.
-Such observations may provide us with significant insights into the structure of spacetime.
-In this context, the LIGO-VIRGO collaboration has been responsible for presenting empirical data on gravitational wave events from the merger of binary black hole system and neutron stars (LIGOScientific:2016aoc, ; LIGOScientific:2017vwq, ).
-In the black hole merger process a ringdown phase arises from the gravitational wave signal, and in this scenario the study of quasinormal modes (QNMs) has been attracting interest in recent years (cardoso2003quasinormal, ; cardoso2019parametrized, ; cruz2020polar, ; anacleto2021quasinormal, ; campos2022quasinormal, ).
-Quasinormal modes are solutions of perturbation equations that have complex characteristic oscillation frequencies that satisfy specific boundary conditions (konoplya2011quasinormal, ).
-The QNMs can be obtained using approximative methods, such as Wentzel-Kramers-Brillouin (WKB) approximation or other numerical methods, in different orders of precision (schutz1985black, ; iyer1987black, ; konoplya2003quasinormal, ; konoplya2019higher, ).
-In higher-dimensional theories, QNMs have gained increasing attention in recent years (Chakraborty:2017qve, ; Dey:2020pth, ; Hashemi:2019jlt, ; konoplya2003quasinormal, ; Cardoso:2003vt, ; Anacleto:2023ntm, ), once one expects such investigations may reveal signals of extra dimensions.
-In turn, recent measurements of the Event Horizon Telescope (EHT) provided images of supermassive black hole shadow at the M87* galaxy center (EventHorizonTelescope:2019dse, ; EventHorizonTelescope:2019ggy, ), in addition to the shadow of compact source Sagittarius A* (Sgr A*), a supermassive black hole in the center of the Milky Way (EventHorizonTelescope:2022xqj, ; EventHorizonTelescope:2022wkp, ).
-It is noteworthy that, in this scenario, the search for physics beyond the SM has been based on analysis of the shadow radius and deviation from the circularity $\Delta C$ of the shadow.
-However, whereas the circularity has been quoted for the M87* measurements, such information for Sgr A* observational data is not available.
+Thus, one expect that black hole physics may provide us with significant insights into the structure of spacetime (Chakraborty:2017qve, ; Dey:2020pth, ; Hashemi:2019jlt, ; Anacleto:2023ntm, ; Cruz:2020emz, ; Cardoso:2003cj, ; Konoplya:2011qq, ; Konoplya:2003ii, ), as well as allow us to test fundamental physics (Yagi:2016jml, ; Goddi:2016qax, ).
+In this context, recent measurements of the Event Horizon Telescope (EHT) provided images of supermassive black hole shadow at the M87* galaxy center (EventHorizonTelescope:2019dse, ; EventHorizonTelescope:2019ggy, ), in addition to the shadow of compact source Sagittarius A* (Sgr A*), a supermassive black hole in the center of the Milky Way (EventHorizonTelescope:2022xqj, ; EventHorizonTelescope:2022wkp, ).
+It is noteworthy that, in this scenario, the search for physics beyond the SM has been based on analysis of the shadow radius and deviation from the circularity $\Delta C$ of the black hole shadow (vagnozzi2022horizon, ).
+However, whereas the limit on the circularity deviation has been quoted for the M87* measurements, imposing $\Delta C\lesssim 10\%$ (EventHorizonTelescope:2019dse, ), such information for Sgr A* is not available.
 Therefore, the approach employed in Ref.
-(Vagnozzi:2019apd, ) cannot be applied to test fundamental physics from the circularity deviations of Sgr A*, and one must recover to investigations based on its shadow size.
-In the present work, we aim to determine conservative constraints on the $\mathrm{AdS_{5}}$ radius of curvature from observational data of the shadow of Sgr A*.
+(Vagnozzi:2019apd, ) cannot be applied to test fundamental physics and to put limits on the $\mathrm{AdS_{5}}$ curvature radius from the circularity deviations of Sgr A* (Vagnozzi2022, ).
+Instead, one must proceed with an investigation based on its shadow size.
+Since the extra dimensions have been proposed in modified frameworks of GR, one expects that they might yield signals on the black hole shadow.
+Therefore, black hole images can be used to test fundamental physics (Yagi:2016jml, ; Vagnozzi2022, ).
+In this work, we perform a study of the shadow radius of Sgr A* aimed at obtaining new and complementary constraints on the extra-dimensional $\mathrm{AdS_{5}}$ curvature radius from observational data.
+In this case, we imposed limits on deviations from the GR’s predictions arising from a braneworld theory.
 For this purpose, we will consider a black hole in the RS background in the weak field regime, described by the Garriga-Tanaka metric (Garriga:1999yh, ).
-This solution has been obtained by assuming a classic approach that provides the linearized brane-world gravity.
-As we will see, our analysis allowed us to place the bounds on the extra-dimensional curvature radius $\ell$ in an alternative way, avoiding the issue of lack of circularity measures for Sgr A* and exploring its inaccurate spin measurements (Banerjee:2022jog, ; Fragione:2020khu, ).
-These findings are supported by the WKB shadow estimate obtained from the study of quasinormal modes in the eikonal limit.
-The organization of this paper is as follows: In section II, we discuss the propagation of a massless scalar field in the background of a black hole on a thin brane-world known as the Randall-Sundrum model (RSII), described by a warped geometry theory.
+This solution has been obtained by assuming a classic approach that provides the linearized braneworld gravity.
+As we will see, our analysis allowed us to place the bounds on the extra-dimensional curvature radius $\ell$ in an alternative way, avoiding the issue of lack of constraints on the deviation from circularity of the shadow measures for Sgr A* and exploring its inaccurate spin measurements (Banerjee:2022jog, ; Fragione:2020khu, ).
+The organization of this paper is as follows: In Sec.
+II, by considering the Randall-Sundrum braneworld scenario, we investigate the null geodesic curves aiming to find the shadow radius of the RSII black hole and discuss the instability of the null circular orbits from the Lyapunov exponent.
 In Sec.
-III, we have determined the quasinormal modes for a scalar perturbation on this background wherein eventual effects of the fifth dimension must become relevant.
-Therefore, we use the WKB approximation, aiming to determine the quasinormal frequencies in the eikonal limit.
-In Sec.
-IV, we investigate the null geodesic curves to find the shadow radius of the RSII black hole, and we compare it with the prediction obtained from the eikonal quasinormal frequencies.
-In this case, we find deviations in the shadow radius of the black hole owing to the extra-dimensional curvature radius.
-From the EHT observational data, we present the imposed constraints on the $\mathrm{AdS_{5}}$ curvature radius $\ell$.
+III, from the EHT observational data of the Sgr A*’s image black hole, we set the imposed constraints on the $\mathrm{AdS_{5}}$ curvature radius $\ell$.
 The concluding remarks and the main results are given in Sec.
-V.
-II Black holes on a thin brane: the radial equation
-In this section, we aim to describe the dynamics of a massless scalar field propagating on a thin brane-world background.
-Thus, we start by considering the Garriga-Tanaka solution that, at the first-order approximation of $GM$, describes the metric of a matter distribution with mass $M$ localized in the brane (Garriga:1999yh, ),
+IV.
+II Corrected Shadow Radius of the Braneworld Black Hole
+In this section, we will study the null geodesic of the Garriga-Tanaka (RSII) black hole, aiming to investigate the influence of the extra dimension on the critical impact and critical radius parameters.
+This approach will also allow us to determine the radius of the shadow of the braneworld black hole.
+From this investigation, we intend to find bounds on the curvature radius of the $\mathrm{AdS_{5}}$ spacetime.
+In this case, we start by considering the Garriga-Tanaka solution that, at the first-order approximation of $GM$, describes the metric of a matter distribution with mass $M$ localized in the brane (Garriga:1999yh, ),
 $ds^{2}=-f(r)dt^{2}+\dfrac{dr^{2}}{g(r)}+r^{2}d\Omega^{2},$ (1)
 where, in the weak field limit $\left(r\gg\ell\right)$ from the horizon of matter after the collapse, we get
-$f(r)=1-\dfrac{2M}{r}-\dfrac{4M\ell^{2}}{3r^{3}},\qquad\text{and}\qquad g(r)=1-\dfrac{2M}{r}-\dfrac{2M\ell^{2}}{3r^{3}},$ (2)
-in which $\ell$ is the curvature radius of $\mathrm{AdS_{5}}$, and we consider unities $c=G=1$.
-In this scenario, we may discuss the propagation of a massless scalar field and obtain the radial equation by using the Klein–Gordon equation
-$\frac{1}{\sqrt{-g}}\partial_{\mu}\left(\sqrt{-g}g^{\mu\nu}\partial_{\nu}\Psi\right)=0.$ (3)
-For the sake of simplicity, one may use the following separation of variables
-$\Psi_{\omega lm}\left(\vec{r},t\right)=\frac{\mathcal{R}_{\omega l}(r)}{r}Y_{l\,m}(\theta,\phi)e^{-i\omega t},$ (4)
-where $Y_{l\,m}(\theta,\phi)$ are the spherical harmonics, $\omega$ is the frequency and $l$ is the multipole number.
-Thus, substituting the expression (4) into equation (3), one finds the following radial equation for $\mathcal{R}_{\omega l}$
-$\Lambda(r)\frac{d}{dr}\left[\Lambda(r)\frac{d\mathcal{R}_{\omega l}}{dr}\right]+\left(\omega^{2}-V_{eff}\right)\mathcal{R}_{wl}=0.$ (5)
-Here, the effective potential is given by
-$V_{eff}=\dfrac{\Lambda(r)}{r}\dfrac{d\Lambda(r)}{dr}+\dfrac{f(r)}{r^{2}}l(l+1),$ (6)
-where $\Lambda(r)=\sqrt{g(r)f(r)}$.
-In order to analyze the radial equation (5), by considering the asymptotic limits, we can introduce a new coordinate $dx=\Lambda(r)^{-1}dr$ (commonly called tortoise coordinate) so that equation (5) reduces to a Schrödinger-type equation, as follows
-$\dfrac{d^{2}\mathcal{R}_{\omega l}}{dx^{2}}+\left(\omega^{2}-V_{eff}\right)\mathcal{R}_{\omega l}=0.$ (7)
-As we will see below, this analysis will allow us to estimate the shadow radius by calculating the quasinormal modes in the eikonal limit, which corresponds to the limit to a large multipole number $l$.
-III Eikonal quasinormal modes in the warped brane-world
-Quasinormal modes are solutions of the perturbation equation (7) that satisfy specific boundary conditions.
-In this case, it is possible to consider two different solutions obtained to the radial equation written in terms of the new coordinate by assuming some set of boundary conditions (berti2009quasinormal, ).
-Furthermore, these solutions present a discrete quasinormal frequency spectrum $\omega=\omega_{R}+i\omega_{I}$, where the real part ($\omega_{R}$) determines the oscillation frequency, while the imaginary part ($\omega_{I}$) determines the decay rate.
-In a regime for which the real part of the QNM frequency is large, one has the called eikonal limit.
-On its turn, WKB approximation has been applied to obtain the quasinormal modes in black hole for decades (schutz1985black, ; iyer1987black, ; konoplya2003quasinormal, ).
-This method is widely employed in the literature due to its practical means of finding approximating solutions compared to numerical methods, in addition to providing accurate predictive results (konoplya2019higher, ; anacleto2021quasinormal, ; campos2022quasinormal, ).
-The WKB method in the eikonal limit yields an accurate estimate for the quasinormal frequencies from the wave equation (7).
-Specifically, the QNM frequencies in this particular limit can be obtained from the following expression
-$\dfrac{i\left(\omega_{n}^{2}-V_{0}\right)}{\sqrt{-2V^{\prime\prime}_{0}}}=n+\dfrac{1}{2},$ (8)
-where $n$ is the harmonic number, $V_{0}$ is the maximum effective potential evaluated at the point $r_{0}$, the prime () denotes derivation with respect to the tortoise coordinate $x$.
-Furthermore, the effective potential (6) at the eikonal limit $(l\rightarrow\infty)$ is $V_{eff}\approx f(r)l^{2}/r^{2}$, so that the equation (8) can be written in the form
-$\omega_{n}\approx\dfrac{l\sqrt{f(r_{0})}}{r_{0}}-\dfrac{i(n+1/2)}{\sqrt{2}}\sqrt{-\dfrac{r_{0}^{2}}{f(r_{0})}\dfrac{d^{2}}{dx^{2}}\left[\dfrac{f(r)}{r^{2}}\right]_{r=r_{0}}},$ (9)
-which corresponds to the quasinormal frequency to the overtone mode $n$.
-In Fig.
-1, one can see that both parts, real and imaginary, of eikonal quasinormal frequencies remain finite.
-In principle, the value of the $\mathrm{AdS_{5}}$ curvature radius $\ell$ must be constrained just from the empirical data such that the quasinormal frequencies will tend to zero in the eikonal limit for large values of the $\ell$-parameter.
-Furthermore, the result (9) obtained by assuming the eikonal limit will be compared with the null geodesics parameters, as we will see in the next section.
-From computing the deviations on the black hole shadow radius in the brane-world model, we will impose bounds on the curvature radius $\mathrm{AdS_{5}}$.
-Figure 1: Real and negative imaginary parts of the eikonal quasinormal frequencies as function of the $\mathrm{AdS_{5}}$ curvature radius.
-IV Corrected Shadow Radius of the Brane-world Black Hole
-In this section, we will study the null geodesic, aiming to investigate the influence of the extra dimension on the critical impact and critical radius parameters.
-This approach will also allow us to determine the radius of the shadow of the Garriga-Tanaka (RSII) black hole.
-From this investigation, we will find bounds on the curvature radius of the $\mathrm{AdS_{5}}$ spacetime.
-We begin by considering the following Lagrangian density
+$f(r)=1-\dfrac{2M}{r}-\dfrac{4M\ell^{2}}{3r^{3}},\quad\text{and}\quad g(r)=1-\dfrac{2M}{r}-\dfrac{2M\ell^{2}}{3r^{3}},$ (2)
+in which $\ell$ is the curvature radius of $\mathrm{AdS_{5}}$, and we consider unities $c=G=1$. We can find the geodesics of the metric (1), considering the following Lagrangian density
 $\mathcal{L}\equiv\frac{1}{2}g_{\mu\nu}\dot{x}^{\mu}\dot{x}^{\nu},$ (3)
@@
 Here, we have that $b=L/E$ is the impact parameter defined as the perpendicular distance (measured from infinity) between the geodesic and the parallel line that passes through the origin.
-Using the following critical conditions, (i) $\mathcal{V}(r_{c})=0$ and (ii) $d\mathcal{V}(r)/dr\big{|}_{r=r_{c}}=0$, we can obtain the critical impact parameter from a system of equations with which one relates the impact parameter to the critical radius, as follows
+Using the following critical conditions, (i) $\mathcal{V}(r_{c})=0$ and (ii) $d\mathcal{V}(r)/dr\big{|}_{r=r_{c}}=0$, we can obtain the critical impact parameter from a system of equations with which one relates the impact parameter to the critical radius, as follows:
 $b_{c}^{2}=\dfrac{r_{c}^{2}}{f(r_{c})}=\dfrac{3r_{c}^{5}}{3r_{c}^{3}-6Mr_{c}^{2}-4M\ell^{2}}.$ (9)
@@
 Specifically, when we analyze the ray in a plane, we observe that any ray of light that starts at a particular angle $\theta$ must keep the same angle throughout its path.
-In Figure 2, we have found some results for the critical impact parameter from the numerical solution of the orbit equations (14) and (15).
+In Fig.
+1, we have found some results for the critical impact parameter from the numerical solution of the orbit Eqs.
+(7) and (8).
 In this case, we have obtained the geodesic trajectory (blue thick curve) computed for the impact parameter $b=5.5M$.
 Therefore, as a direct influence of the existence of the extra-dimensional parameter $\ell$, we observed an increase in the absorption section.
-Furthermore, in Figure 2, whereas the black disk depicts the limit of the event horizon, the critical radius for the photon sphere is shown as an internal dotted circle.
+Furthermore, in Fig.
+1, whereas the black disk depicts the limit of the event horizon, the critical radius for the photon sphere is shown as an internal dotted circle.
 In turn, the critical impact parameter (black hole shadow) is shown by an external dashed circle.
 Figure 1: Null geodesic in polar coordinates for different values of $\ell$, in which we consider a test beam of light by keeping the impact parameter, $b=5.5M$, fixed.
+In this scenario, it is important to verify the instability of null circular orbits.
+For this, we have two important quantities, namely the Lyapunov exponent $\lambda=\sqrt{\mathcal{V}^{\prime\prime}(r)/2\dot{t}^{2}}$ that determines the unstable timescale and the angular velocity $\Omega_{c}=\dot{\phi}/\dot{t}$ for the last null circular orbit (cardoso2009geodesic, ; Anacleto:2021qoe, ; Campos:2021sff, ).
+The instability exponent is defined by the ratio of the Lyapunov exponent to the angular velocity.
+In Fig.
+2, we see the variation of the instability exponent in relation to the radius $r_{c}$.
+The maximum of the instability increases with the increase in the contribution coming from the extra dimension.
+On the other hand, the instability decreases as $r_{c}$ increases.
+It is worth mentioning that this result is directly sensitive to the geometry, so this analysis is expected to eventually be used to distinguish between Schwarzschild and braneworld black holes.
+Figure 2: Lyapunov exponent normalized by the angular velocity $\left(\lambda/\Omega_{c}\right)$, as function of $r_{c}/M$ for some values of $\ell$.
 Finally, one can estimate the size of the radius of the black hole shadow by using the study of geodesic curves that we might express via celestial coordinates (vazquez2003strong, )
@@
 $R_{s}\equiv\sqrt{\alpha^{2}+\beta^{2}}=\sqrt{\dfrac{3r_{c}^{5}}{3r_{c}^{3}-6Mr_{c}^{2}-4M\ell^{2}}}.$ (13)
-Let us now expand the radius of the shadow up to order $\left(\ell^{2}/M^{2}\right)$ by assuming that the contribution arising from extra-dimension is a small correction to $R_{s}$, $\ell/M\ll 1$, so that
+Let us now expand the radius of the shadow up to order $\left(\ell^{2}/M^{2}\right)$ by assuming that the contribution arising from extra dimension is a small correction to $R_{s}$, i.e., $\ell/M\ll 1$, so that
 $R_{s}\approx 3\sqrt{3}M\left[1+\frac{2}{27}\frac{\ell^{2}}{M^{2}}+\mathcal{O}\left(\frac{\ell^{4}}{M^{4}}\right)\right].$ (14)
-On the other hand, for a spherically symmetric and asymptotically flat static geometry, Cardoso et al.
-(cardoso2009geodesic, ) showed that, in the eikonal limit, the quasinormal modes are determined by the parameters of the null circular geodesic so that $\Omega_{c}=\dot{\phi}/\dot{t}$ is the angular velocity for the last null circular orbit, and $\lambda=\sqrt{\mathcal{V}^{\prime\prime}(r)/2\dot{t}^{2}}$ is the Lyapunov exponent that determines the unstable time scale of the orbit.
-By directly comparing with equation (9), we can write the equation for quasinormal frequencies in the form
-$\omega_{n}=\Omega_{c}l-i(n+1/2)|\lambda|.$ (22)
-It is worth to mention that there are violations in this relationship, as shown by Konoplya and Stuchlík (konoplya2017eikonal, ) when considering a black hole in the Einstein-Loverlock theory.
-Since we know the form of the angular velocity $\Omega_{c}$, one can use equations (12) and (16) to obtain its relationship with the critical impact parameter $\Omega_{c}=\left(f(r_{c})b_{c}\right)/r_{c}^{2}=1/b_{c}$.
-This equation highlights the correspondence relationship between the maximum value point of the effective potential ($r_{0}$) and the radius of the photon circular orbit ($r_{c}$) [see equations (9) and (22)].
-Finally, in the eikonal regime $l\gg 1$, we have the relationship between the real part of the quasinormal modes and the shadow radius,
-$Re(\omega)=\lim_{l\gg 1}\left(\dfrac{l}{R_{s}}\right).$ (23)
-Notice that the estimate (23) may be employed to confirm the prediction (21).
-Hence, we can use it to compare with experimental data in order to constrain the found deviations.
-Besides, equation (22) shows that the real and imaginary parts of the quasinormal modes in the eikonal limit are related to the orbital angular velocity and the instability time scale of circular null geodesics, respectively.
-The instability of null circular orbits can be determined by the instability exponent, which is defined by the ratio of the Lyapunov exponent to the angular frequency.
-In Fig.3, we see the variation of the instability exponent in relation to the radius $r_{c}$.
-The maximum of the instability increases with the increase in the contribution coming from the extra dimension.
-On the other hand, the instability decreases as $r_{c}$ increases.
-Figure 3: Lyapunov exponent normalized by the angular velocity $\left(\lambda/\Omega_{c}\right)$, as function of $r_{c}/M$ for some values of $\ell$.
-IV.1 Constraints on the $\mathrm{AdS_{5}}$ curvature radius
+Notice that the extra-dimensional deviation can be understood as a correction to the Schwarzschild’s shadow radius estimate.
+In the next section, we will find limits on the correction to the shadow radius.
+III Constraints on the $\mathrm{AdS_{5}}$ curvature radius
 As we saw above, we obtained deviations in the black hole shadow radius owing to the existence of an extra dimension.
@@
 Although this independent approach allows us to search for traces of extra dimensions, imposing constraints on deviations found, once the curvature radius $\ell$ is positively correlated to the shadow radius, and given the low-precision measurements, thus we must have a weakening in the predicted bounds compared to sub-mm gravity experiments, for instance.
-In our analysis, we will follow the methodology developed in (vagnozzi2022horizon, ), aiming to obtain bounds on the $\mathrm{AdS_{5}}$ curvature radius.
-Therefore, the observed angular radius – which has ring-like features in the horizon-scale EHT image of Sgr A* –, must be compared to the theoretically determined angular radius of the black hole shadow.
+In our analysis, we aim to obtain bounds on the $\mathrm{AdS_{5}}$ curvature radius.
+Therefore, the observed angular radius—which has ringlike features in the horizon-scale EHT image of Sgr A*—must be compared to the theoretically determined angular radius of the black hole shadow.
 From the precise determination of both mass and distance of Sgr A*, the central supermassive black hole of our galaxy, we can verify the accordance, within the uncertainty allowed, between the predicted theoretical results by this study and the empirical data found by the EHT observations (akiyama2022millimeter, ).
-On its turn, the mass and distance of Sgr A* have been studied over the last few decades by exploring the dynamics of star clusters and, in particular, the movement of stars close to the center of the galaxy.
+On its turn, the mass and distance of Sgr A* have been studied over the past few decades by exploring the dynamics of star clusters and, in particular, the movement of stars close to the center of the galaxy.
 According to Ref.
-(vagnozzi2022horizon, ), the shadow radius measurements are restricted to the following regions determined by the average between the estimates based on the VLTI and Keck instruments, within $1\sigma$ and $2\sigma$ intervals, respectively,
-$4.55\lesssim R_{s}/M\lesssim 5.22,\qquad\text{and}\qquad 4.21\lesssim R_{s}/M\lesssim 5.56$ (24)
-In Figure 4a, we show the expected deviations in the behavior of the shadow radius owing to the extra dimension.
+(EventHorizonTelescope:2022xqj, ), the shadow radius measurements are restricted to the following region determined by the estimates based on the Keck and Very Large Telescope Interferometer (VLTI) instruments, within $1\sigma$ confidence interval, respectively,
+$4.5\lesssim R_{s}/M\lesssim 5.5,\quad\text{and}\quad 4.3\lesssim R_{s}/M\lesssim 5.3$ (15)
+In Fig.
+3a, we show the expected deviations in the behavior of the shadow radius owing to the extra dimension.
 As we see, the increase in the $\mathrm{AdS_{5}}$ curvature radius $\ell$ results in a consequent increase in the radius of the black hole shadow.
-In our analysis, we have considered the average in the mass from KECK- and VLTI-based estimates.
-Therefore, from the empirical data obtained by observations of EHT, we can impose the limits on curvature radius at the $95\%$ confidence level (CL), such that $\ell\lesssim 6.5\times 10^{9}\,\mathrm{m}\simeq 4.3\times 10^{-2}\,\mathrm{AU}$, as shown in Figure 4b.
+In our analysis, we have considered the mass from Keck- and VLTI-based estimates (vagnozzi2022horizon, ).
+Therefore, from the empirical data obtained by observations of EHT, we can impose the upper limits on curvature radius at the $68\%$ confidence level (CL), such that
+$\ell_{\left(\text{Keck}\right)}\lesssim 5.7\times 10^{9}\,\mathrm{m}\simeq 3.8\times 10^{-2}\,\mathrm{AU},\quad\text{and}\quad\ell_{\left(\text{VLTI}\right)}\lesssim 2.4\times 10^{9}\,\mathrm{m}\simeq 1.6\times 10^{-2}\,\mathrm{AU},$ (16)
+where $1\mathrm{AU}$ is one astronomical unit. These results are shown in Fig. 3b.
 At this point, we can compare our results with another which investigates similar physical systems.
 In this case, the absence of deviations on observation of the highly circular shadow of black hole M87* provides constraints on the $\mathrm{AdS_{5}}$ radius so that $\ell\lesssim 170\,\mathrm{AU}$ (Vagnozzi:2019apd, ).
-In turn, the first bounds obtained on the extra dimensions by the study of the time-lag between the detection of gravitational waves (GW) and electromagnetic (EM) signals from the merging of a binary system detected by the event GW170817 have established at $68\%$ CL an upper limit of $\ell\lesssim 0.535\,\mathrm{Mpc}\simeq 1.10\times 10^{11}\,\mathrm{AU}$ (Visinelli:2017bny, ), whereas our analysis found $\ell\lesssim 1.0\times 10^{-2}\,\mathrm{AU}$ for the same confidence interval.
-It is worth mentioning that, as we have seen, our study yields stronger constraints than those found by similar physical systems, according to the best of our knowledge.
-Ultimately, although we obtained weaker constraints compared to results obtained from sub-mm scale gravity experiments, this approach has set independent limits and can aid the investigation of the geometric structure of the universe.
-Figure 4: The left panel shows the changes in the behavior of the shadow radius with the variation of the $\mathrm{AdS_{5}}$ radius $\ell$.
-In turn, in the right panel, we found bounds from the direct comparison between the results obtained by our theoretical analysis with the EHT experimental data.
-V Final Remarks
+In turn, the first bounds obtained on the extra dimensions by the study of the time lag between the detection of gravitational waves (GW) and electromagnetic (EM) signals from the merging of a binary system detected by the event GW170817 have established at $68\%$ CL an upper limit of $\ell\lesssim 0.535\,\mathrm{Mpc}\simeq 1.10\times 10^{11}\,\mathrm{AU}$ (Visinelli:2017bny, ), whereas our analysis found $\ell_{\left(\text{Keck}\right)}\lesssim 3.8\times 10^{-2}\,\mathrm{AU}\left(\ell_{\left(\text{VLTI}\right)}\lesssim 1.6\times 10^{-2}\,\mathrm{AU}\right)$ for the same confidence interval.
+It is worth mentioning that, as we have seen, our study yields more stringent constraints than those found by similar physical systems, according to the best of our knowledge.
+Finally, although we obtained weaker constraints compared to results obtained from sub-mm scale gravity experiments, this approach has set independent limits and can aid the investigation of the geometric structure of the Universe.
+Figure 3: (a) Shows the changes in the behavior of the shadow radius with the variation of the $\mathrm{AdS_{5}}$ radius $\ell$.
+In turn, (b) we found bounds from the direct comparison between the results obtained by our theoretical analysis with the EHT experimental data.
+IV Final Remarks
 Recently, higher-dimensional theories have drawn increasing attention.
 As has been pointed out, the seek for evidence of large extra dimensions has been performed in several experimental researches, making its detection a challenging problem in itself.
-In this work, we have studied the quasinormal modes of a black hole in a brane-world model, aiming to search for traces of extra dimensions.
-By employing the WKB method, we compute the eikonal quasinormal frequencies.
-This complementary analysis also allowed us to confirm the theoretical prediction for the black hole shadow radius from the convergence between the WKB- and null geodesics-method estimates.
-From recent measurements of the shadow image of Sgr A* – a supermassive black hole in the center of our galaxy – we estimate deviations owing to an extra dimension in the RSII scenario on the theoretical shadow radius predicted.
-Given the inaccuracy involving the measured spin of Sgr A*, which may even be zero but still can provide compatible results with recent measures (Fragione:2020khu, ), we explore the possibility of describing them as a non-rotating RSII black hole.
-In this case, we have considered the Garriga-Tanaka solution that takes a classical approach to linearized brane-world gravity and describes a black hole solution in the weak field limit.
+By investigating the null geodesic curves, we find theoretical prediction for the black hole shadow radius, aiming to search for traces of extra dimensions.
+From recent measurements of the shadow image of Sgr A*—a supermassive black hole in the center of our Galaxy—we estimate deviations owing to an extra dimension in the RSII scenario on the theoretical shadow radius predicted.
+Given the inaccuracy involving the measured spin of Sgr A*, which may even be zero but still can provide compatible results with recent measures (Fragione:2020khu, ), we explore the possibility of describing them as a nonrotating RSII black hole.
+In this case, we have considered the Garriga-Tanaka solution that takes a classical approach to linearized braneworld gravity and describes a black hole solution in the weak field limit.
 As we saw, the curvature radius of the $\mathrm{AdS_{5}}$ spacetime is positively correlated to the shadow radius so that the increasing/decreasing in one leads to correspondingly increasing/decreasing in the other.
 Furthermore, we found that $\ell=0$, i.e., in the absence of new physics coming from extra dimension, one recovers the standard estimates expected for the Schwarzschild background.
-By investigating the null geodesic curves, we determine the shadow of the RSII black hole.
-From recent observational data of the shadow of Sgr A* released by the Event Horizon Telescope collaboration, one may impose constraints on the curvature radius of $\mathrm{AdS_{5}}$ spacetime.
-Therefore, we find more stringent limits on $\ell\lesssim 4.3\times 10^{-2}\,\mathrm{AU}$ (at $95\%$ CL) than previously obtained for similar physical systems, to the best of our knowledge.
-Moreover, non-zero predicted curvature radius also leads to agreement with the observed shadow radius within $68\%$ confidence interval, providing $\ell\lesssim 1.0\times 10^{-2}\,\mathrm{AU}$.
-We would like to highlight that although our analysis provides weaker bounds than those obtained from collider experiments or sub-mm precision gravity tests, which have set limits on the non-Newtonian force, in this study, we have provided conservative bounds since limits ourselves to investigate the shadow radius in a non-rotating RSII-type black hole background.
-Further, the bounds reported by us have been obtained in a manner complementary to the studies previously mentioned.
-Finally, the newly begun era of observational astronomy opens several avenues for future research that might be explored to obtain new constraints on the $\mathrm{AdS_{5}}$ curvature radius, for instance, or even to test extensions of general relativity, therefore shedding light into seeking for physics beyond the SM.
+From recent observational data of the shadow of Sgr A* released by the EHT collaboration, one may impose constraints on the curvature radius of $\mathrm{AdS_{5}}$ spacetime.
+Therefore, by considering the Keck and VLTI estimates, we find more stringent limits on the $\mathrm{AdS_{5}}$ curvature radius, $\ell_{\left(\text{Keck}\right)}\lesssim 3.8\times 10^{-2}\,\mathrm{AU}$ and $\ell_{\left(\text{VLTI}\right)}\lesssim 1.6\times 10^{-2}\,\mathrm{AU}$ (at $68\%$ CL), than previously obtained for similar physical systems, to the best of our knowledge.
+Besides, a nonzero predicted curvature radius leads to agreement with the observed shadow radius at the $68\%$ confidence interval.
+We would like to highlight that, although our analysis provides weaker bounds than those obtained from collider experiments or sub-mm precision gravity tests, which have set limits on the non-Newtonian force, in this study, we have provided conservative bounds since we investigated the shadow radius in a nonrotating RSII-type black hole background.
+Further, the bounds we reported were obtained independently and in a manner complementary to the previously mentioned studies.
+Finally, the newly begun era of observational astronomy opens several avenues for future research that might be explored for testing fundamental physics—and so getting new constraints on the $\mathrm{AdS_{5}}$ curvature radius, for instance—or even searching eventual deviations on GR, shedding light into seeking physics beyond the SM.
 Acknowledgements.
-We would like to thank CNPq, CAPES and CNPq/PRONEX/FAPESQ-PB (Grant No.
+We thank the referee for the valuable comments and suggestions that have enhanced this work.
+We would like to thank CNPq, CAPES, and CNPq/PRONEX/FAPESQ-PB (Grant No.
 $165/2018$), for partial financial support.
-FAB acknowledges support from CNPq (Grant No.
+F.A.B.
+acknowledges support from CNPq (Grant No.
 309092/2022-1).
-ASL acknowledges support from CAPES (Grant No.
+A.S.L.
+acknowledges support from CAPES (Grant No.
 88887.800922/2023-00).
-JAVC would like to thank FAPESQ-PB/CNPq (Grant No.
+J.A.V.C.
+would like to thank FAPESQ-PB/CNPq (Grant No.
 077/2022) for financial support.